An elevator motor in a high-rise building can do 3500 kJ of work in 5 min. Find the power developed by

the motor.


1
Expert's answer
2021-05-20T18:24:28-0400
"P=\dfrac{W}{t}=\dfrac{3.5\cdot10^6\ J}{5\cdot60\ s}=11667\ W=11.6\ kW."